# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v1

@@ Abstract.
This paper introduces PAT, a prefix-aware attention kernel implementation for LLM decoding that organizes execution with a pack-forward-merge paradigm. PAT packs queries by shared prefix to reduce repeated memory accesses, runs a customized multi-tile kernel to achieve high resource efficiency. It further applies practical multi-stream forwarding and KV splitting to reduce resource bubbles. The final merge performs online softmax with negligible overhead. We implement PAT as an off-the-shelf plugin for vLLM. Evaluation on both real-world and synthetic workloads shows that PAT reduces attention latency by 67.4% on average and TPOT by 13.6-83.4% under the same configurations against state-of-the-art attention kernels.
@@ 1. Introduction
Tranformer-based Large Language Models (LLMs) have rapidly advanced across diverse domains, including conversational assistants (Dam et al., 2024; Dong et al., 2023), code generation (Anthropic, 2025; Hui et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020; Yao et al., 2025), and agent/tool workflows (Microsoft, 2025; Yuan et al., 2024). As deployment grows, latency and resource efficiency of online inference have become increasingly critical. While prior work has explored optimizations in memory management (Kwon et al., 2023; Xu et al., 2025), scheduling (Zhong et al., 2024; Sun et al., 2024), quantization (Lin et al., 2024b; Chee et al., 2023), and architecture (Zhou et al., 2022; Liu et al., 2024), there remain new challenges and opportunities. We highlight two emerging trends below.
Longer contexts and outputs. Context length has scaled to millions of tokens (AI, 2025), while techniques like Chain-of-Thought (CoT) (Wei et al., 2022) demand longer outputs. As a result, decode attention (i.e., attention operations in the decoding stage) increasingly dominates end-to-end latency due to repeated loading of growing KV caches from global to on-chip memory, which poses a memory-bound challenge.
Shared prefixes across requests. System prompts, RAG documents, and agent templates introduce multi-level shared prefixes across requests, creating an optimization opportunity. Existing systems (Zheng et al., 2024b; Kwon et al., 2023) implement prefix KV cache reuse, which reduces memory footprint by storing and reusing KV cache across requests. However, it can not reduce global memory accesses, which is the bottleneck for decode attention, thereby leading to higher attention latency.
Recent attention kernel implementations aim to reduce attention latency (Shah et al., 2024; Ye et al., 2025, 2024b; Pan et al., 2025; Song et al., 2024; Zhu et al., 2024), but still face redundant KV cache loads from global memory and inefficient resource utilization. Query-centric kernels (Dao et al., 2022; Dao, 2023; Ye et al., 2025; Song et al., 2024) follow a one-query-per-CTA strategy to map each query and corresponding KV cache into an independent Cooperative Thread Array (CTA; also known as a thread block) for execution on the GPU, causing redundant global memory loads for shared prefixes. KV-centric kernels (Pan et al., 2025; Zhu et al., 2024; Yao et al., 2024; Ye et al., 2024b) instead pack the KV cache of cross-query shared prefix into one CTA to reduce redundant memory accesses, but adopt a one-size-fits-all design that fixes the tile size of the GPU kernel (NVIDIA, 2025a) and applies padding to fill the tile, which wastes on-chip memory and limits CTA concurrency.
@@
To address these challenges, we present PAT, an attention kernel implementation for LLM decode stage with a pack–forward–merge execution paradigm. Specifically, in the pack stage5), PAT employs a prefix-aware pack scheduler and a lazy update mechanism, which efficiently pack decode batches into CTAs to reduce redundant global memory accesses with negligible overhead. It further designs a multi-tile kernel based on resource-efficiency analysis, so as to adapt to CTA dynamicity and avoid on-chip memory waste. In the forward stage6), PAT designs a multi-stream forward and long-KV split strategy, which parallelizes the multi-tile kernel across multiple CUDA streams and splits CTAs with excessively long KV lengths. These designs enable PAT to eliminate execution bubbles and achieve high global memory bandwidth utilization. In the merge stage7), PAT applies a lightweight kernel with online softmax (Dao et al., 2022), which merges partial results across CTAs for each query.
@@
    We evaluate PAT on synthetic and real-world workloads, demonstrating that compared with state-of-the-art baselines, it reduces attention latency by 67.4% on average under the same decode batches (§8.3) and lowers TPOT by 13.6–83.4% under the same request rate (§8.4).
@@ 2.2. Memory Bottleneck of Attention
Driven by the inference scaling law (Wu et al., 2025), LLMs are trending toward longer contexts and longer outputs. For instance, Llama-4 supports up to 10 million input tokens (AI, 2025), and Chain-of-Thought prompting (Wei et al., 2022) has significantly increased output lengths for complex reasoning.
Although KV Caching avoids repeated computation, it has shifted the bottleneck from compute to memory access: each decode step must fetch a growing amount of cached KK/$V$ data from global memory to on-chip memory (Zadouri et al., 2025; Behnam et al., 2025). As input and output lengths grow, this access cost dominates inference latency. For models like Llama-3-8B and Qwen3-8B, decode attention can contribute up to 68% of the total latency, as shown in Figure 1.
@@ 2.3. GPU Execution Model
Table 1. Memory hierarchy of A100-SXM4-80GB (NVIDIA, 2020; Abdelkhalik et al., 2022).
@@
When executing an attention kernel, the workload is divided into CTAs, scheduled across SMs. During this process, the required KV cache data is transferred from slow global memory up the hierarchy into fast on-chip shared memory, where it is finally loaded into registers for computation. State-of-the-art implementations (Dao et al., 2022; Dao, 2023; Ye et al., 2025) are designed to exploit this memory hierarchy. For example, FlashAttention (Dao et al., 2022; NVIDIA, 2025a) partitions K/V caches into small tiles along the sequence length (Figure 3a) and processes them in a pipelined fashion: while one tile is computed, the next is prefetched asynchronously (Figure 3b), reducing memory latency.
@@ 3.1. Shared Prefixes and Prefix Reuse
Shared prefixes are prevalent in modern LLM workloads (ope, 2025; Wang et al., 2020; Jin et al., 2024), with studies showing that 40–62% of KV cache prefixes are shared across requests (Qin et al., 2025; Wang et al., 2025). These prefixes often form multi-level hierarchies (Zheng et al., 2024b; Pan et al., 2025). For example, a base system prompt shared by all users, followed by task-specific prefixes shared within subgroups.
Systems like SGLang(Zheng et al., 2024b) and vLLM(Kwon et al., 2023) support prefix reuse by mapping shared logical prefixes to a single physical copy in GPU global memory. This reduces KV cache memory usage and improves concurrency. However, it can not reduce global memory access, which remains the decode attention bottleneck. To understand why, we analyze two key inefficiencies in current attention kernels under shared-prefix scenarios: redundant memory access (§3.2) and underutilized hardware resources (§3.3) as follows.
@@ 3.2. Redundant Memory Access
A major inefficiency in current attention kernels stems from redundant memory accesses. To execute a decode batch, attention kernels use a packing strategy that groups queries and their KV caches into CTAs. In the decode batch with 4 queries shown in Figure 4a, existing kernels’ query-centric paradigm adopts a one-query-per-CTA packing strategy (Shah et al., 2024; Ye et al., 2025), where each query and its KV blocks are independently assigned to a CTA (Figure 4b). While simple to schedule, this strategy causes shared KV prefixes (e.g., KV-0, KV-1) to be repeatedly loaded from slow global memory into on-chip memory. Although L2 cache offers partial reuse, it is limited by size and bandwidth.
@@ 3.3. Two-Dimension Resource Inefficiencies
Furthermore, existing attention kernels exhibit inefficient hardware utilization. Both query-centric and KV-centric designs use a tiled pipeline (Figure 3) to overlap memory access and computation. In this paradigm, selecting appropriate tile sizes (i.e., query tile size $m$ and KV tile size $n$ from Figure 3b) is critical for resource efficiency (Dao et al., 2022). However, existing kernels adopt a one-size-fits-all design, which specifies a single, hard-coded tile size for all CTAs (e.g., $m=64,n=32$ (Shah et al., 2024)).
As shown in Figure 4b-c, this static approach ignores the dynamic nature of LLM workloads, leading to significant resource inefficiencies along two dimensions [^1] While specific prior works (Pan et al., 2025) design one more query tile sizes $m$, it is still insufficient for handling highly dynamic query numbers and KV lengths. in Figure 5(b): (1) Memory Waste ($\mathbb{I}_{mem}$): When fewer than $m$ queries share a KV prefix, CTAs must pad inputs, wasting shared memory and registers to store unused data. (2) Execution Bubble ($\mathbb{I}_{exe}$): Due to varying KV lengths across CTAs, fixed-size tiling causes imbalanced workloads, leaving SMs underutilized in the tail stages of execution.
@@ 3.4. Insight: Pack-forward-merge Paradigm
Building on the observations above, we identify two design principles for efficient decode attention: (1) Intra-CTA KV Cache Sharing: pack queries with shared KV prefixes into the same CTA (Figure 4c) to enable KV reuse in shared memory and avoid redundant global memory access. (2) Resource-Efficient Kernel Design: tailor kernel implementations to GPU architecture and CTA configurations to sustain high memory bandwidth and minimize resource inefficiencies.
@@ 4. Overview
To address these challenges, we design PAT, an attention kernel implementation that follows the pack-forward-merge paradigm. It serves as a backend for the serving system vLLM (Kwon et al., 2023). In the pack stage, PAT adopts a profit-model-based heuristic packing strategy to aggregate queries that share KV into a CTA, so as to mitigate redundant global memory accesses (§5.1). It further designs multi-tile kernels and a runtime tile selector to choose an efficient kernel for each CTA (§5.2). In the forward stage, PAT adopts multi-stream execution and a long-KV split strategy to enable efficient kernel execution, which reduces execution bubbles (§6). Finally, in the merge stage, PAT uses a lightweight kernel based on online softmax to merge each query’s intermediate results across CTAs and produce the final output (§7).
@@ 5.1. Pack Scheduler
Problem Formulation Since decode attention is memory-bound (§2.3 and §3.2), the pack stage aims to minimize global memory accesses for a given decode batch (Figure 6a). The input is a batch of queries plus a block table where each row lists block IDs for a query; a shared prefix appears as identical leading block IDs across rows. The output is a partition $\mathcal{P}=\{P_{1},P_{2},\dots\}$ of CTAs, where each $P_{i}$ packs queries sharing successive identical prefix blocks (queries may be split across CTAs). We seek $\mathcal{P}^{\ast}$ that minimizes memory accesses, comprised of loaded KV cache size and per-query intermediate reads/writes due to splits/merges. The search space grows exponentially with query count and prefix lengths  (Pan et al., 2025), so an exact solver is impractical for online serving, and we turn to a heuristic pack scheduler described next.
@@
In practice, the length of shared KV $l_{u}\geq 16$ since sharing is performed at the granularity of KV blocks (Kwon et al., 2023; Shah et al., 2024), whose sizes are typically larger than 16. Therefore, packing a node into a CTA yields a positive profit.
@@
Lazy Update. Although the pack scheduler is linear, its overhead can grow with batch size and number of KV blocks. We mitigate this with a lazy-update strategy that (1) reuses a scheduling result across continuous-batching iterations until the block table changes (e.g., request arrivals/departures or new KV block assignments) and (2) moves the scheduler into the serving system and runs it asynchronously once the block table is available. These reduce scheduler invocations from once per transformer layer to once several continuous-batching iterations and overlap scheduling with pre-attention task (e.g., LayerNorm, QKV projection), thereby substantially reducing the exposed scheduling latency (see Appendix B) without affecting model accuracy.
@@ 5.2. Multi-tile Kernel
Insight and Approach. Given packed CTAs, PAT must choose per-CTA Q-tile $m$ and KV-tile $n$ to maximize resource efficiency. Because query sizes and KV lengths vary widely, single–tile-size kernels with padding (e.g., prior works (Dao, 2023; Ye et al., 2025; Pan et al., 2025; Zhu et al., 2024)) cannot adapt and suffer poor utilization (§3.3). To better exploit GPU shared memory and registers, we introduce a customized tiled attention kernel design that consists of: (1) a multi-tile kernel suite, where feasible $(m,n)$ configurations are derived from offline hardware and CTA-constraint analysis and implemented as resource-efficient kernels; and (2) a tile-size selector, an online decision-tree that selects the per-CTA $(m,n)$ to balance performance and parallelism.
@@ 1 $\mathcal{P}\leftarrow\emptyset$;
①Register and shared-memory constraints (upper bounds on $m,n$). Increasing $m$ or $n$ raises CTA shared-memory and register usage and can induce register spilling (NVIDIA, 2025b). To keep the kernel within hardware limits, we enforce two bounds: (1) Shared-memory constraint. One CTA’s shared-memory usage comprises the Q tile, K/V tile, and intermediate results (data type size $b^{\prime}$, usually higher precision). It must not exceed per-SM shared memory $S_{\mathrm{smem}}$:
@@
③CUTLASS constraint (lower bounds for $m$ and $n$): Efficient use of CUTLASS/CuTe MMA requires both tile sizes to be powers of two and at least [^5] It depends on data format; e.g., the minimum tile size is 32 for int8. 16 (NVIDIA, 2025c):
@@
Kernel equivalence. Per the offline constraint solving above, the candidate $(m,n)$ configurations sustain high bandwidth utilization, and decode-attention latency is dominated by global memory bandwidth. Thus, for decode batches without shared prefixes or execution bubbles, these configurations are performance-equivalent in bandwidth and latency. To validate this, we executed a decode batch without prefixes (KV length $1024$) under various configurations and used batch size 1134113466611341134 is a common multiple of CTA concurrency across configurations on A100, avoiding execution bubbles.. As shown in Figure 7(c) and Figure 7(d), all candidate configurations (underlined) sustain $83\text{–}86\%$ bandwidth utilization and exhibit similar end-to-end latency (difference $2\%$). This demonstrates that, in the absence of prefixes and bubbles, varying the tile configuration within the feasible set does not change CTA performance.

### v2

@@ Abstract.
This paper introduces PAT, a prefix-aware attention kernel implementation for LLM decoding that organizes execution with a pack-forward-merge paradigm. PAT packs queries by shared prefix to reduce repeated memory accesses, runs a customized multi-tile kernel to achieve high resource efficiency. It further applies practical multi-stream forwarding and KV splitting to reduce resource bubbles. The final merge performs online softmax with negligible overhead. We implement PAT as an off-the-shelf plugin for vLLM. Evaluation on both real-world and synthetic workloads shows that PAT reduces attention latency by 53.5% on average and TPOT by 17.0-93.1% under the same configurations against state-of-the-art attention kernels. PAT’s source code is publicly available at https://github.com/flashserve/PAT.
Large Language Models, LLM Inference, Prefix Aware Attention, GPU Kernel Scheduling
copyright: rightsretainedjournalyear: 2026conference: 31th ACM International Conference on Architectural Support for Programming Languages and Operating Systems; March 22–March 26, 2026; Pittsburgh, USAbooktitle: Proceedings of the 31th ACM International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS ’26), March 32–March 26, 2026, Pittsburgh, USAdoi: 10.0000/0000000.0000000isbn: 000-0-0000-0000-0/00/00ccs: Computing methodologies Machine learningccs: Computer systems organization Cloud computing
@@ 1. Introduction
Tranformer-based Large Language Models (LLMs) have rapidly advanced across diverse domains, including conversational assistants (dam2024complete; dong2023towards), code generation (anthropic2025claudecode; hui2024qwencoder), retrieval-augmented generation (RAG) (lewis2020retrieval; yao2025cacheblend), and agent/tool workflows (microsoft2025autogen; yuan2024easytool). As deployment grows, latency and resource efficiency of online inference have become increasingly critical. While prior work has explored optimizations in memory management (kwon2023pagedattention; xu2025ellm), scheduling (zhong2024distserve; sun2024llumnix), quantization (lin2024awq; chee2023quip), and architecture (zhou2022mixture; liu2024deepseek), there remain new challenges and opportunities. We highlight two emerging trends below.
Longer contexts and outputs. Context length has scaled to millions of tokens (meta2025llama4), while techniques like Chain-of-Thought (CoT) (wei2022chain) demand longer outputs. As a result, decode attention (i.e., attention operations in the decoding stage) increasingly dominates end-to-end latency due to repeated loading of growing KV caches from global to on-chip memory, which poses a memory-bound challenge.
Shared prefixes across requests. System prompts, RAG documents, and agent templates introduce multi-level shared prefixes across requests, creating an optimization opportunity. Existing systems (zheng2024sglang; kwon2023pagedattention) implement prefix KV cache reuse, which reduces memory footprint by storing and reusing KV cache across requests. However, it can not reduce global memory accesses, which is the bottleneck for decode attention, thereby leading to higher attention latency.
Recent attention kernel implementations aim to reduce attention latency (shah2024flashattention3; ye2025flashinfer; ye2024chunkattention; pan2025fasttree; song2024tackling; zhu2024relayattention), but still face redundant KV cache loads from global memory and inefficient resource utilization. Query-centric kernels (dao2022flashattention; dao2023flashattention2; ye2025flashinfer; song2024tackling) follow a one-query-per-CTA strategy to map each query and corresponding KV cache into an independent Cooperative Thread Array (CTA; also known as a thread block) for execution on the GPU, causing redundant global memory loads for shared prefixes. KV-centric kernels (pan2025fasttree; zhu2024relayattention; yao2024deft; ye2024chunkattention) instead pack the KV cache of cross-query shared prefix into one CTA to reduce redundant memory accesses, but adopt a one-size-fits-all design that fixes the tile size of the GPU kernel (nvidia2025cudabestpractices) and applies padding to fill the tile, which wastes on-chip memory and limits CTA concurrency.
@@
To address these challenges, we present PAT, an attention kernel implementation for LLM decode stage with a pack–forward–merge execution paradigm. Specifically, in the pack stage5), PAT employs a prefix-aware pack scheduler and a lazy update mechanism, which efficiently pack decode batches into CTAs to reduce redundant global memory accesses with negligible overhead. It further designs a multi-tile kernel based on resource-efficiency analysis, so as to adapt to CTA dynamicity and avoid on-chip memory waste. In the forward stage6), PAT designs a multi-stream forward and long-KV split strategy, which parallelizes the multi-tile kernel across multiple CUDA streams and splits CTAs with excessively long KV lengths. These designs enable PAT to eliminate execution bubbles and achieve high global memory bandwidth utilization. In the merge stage7), PAT applies a lightweight kernel with online softmax (dao2022flashattention), which merges partial results across CTAs for each query.
@@
    We evaluate PAT on synthetic and real-world workloads, demonstrating that compared with state-of-the-art baselines, it reduces attention latency by 67.4% on average under the same decode batches (§8.3) and lowers TPOT by 17.0-93.1% under the same request rate (§8.4).
@@ 2.2. Memory Bottleneck of Attention
Driven by the inference scaling law (wu2025inference), LLMs are trending toward longer contexts and longer outputs. For instance, Llama-4 supports up to 10 million input tokens (meta2025llama4), and Chain-of-Thought prompting (wei2022chain) has significantly increased output lengths for complex reasoning.
Although KV Caching avoids repeated computation, it has shifted the bottleneck from compute to memory access: each decode step must fetch a growing amount of cached KK/$V$ data from global memory to on-chip memory (zadouri2025hardware; behnam2025rocketkv). As input and output lengths grow, this access cost dominates inference latency. For models like Llama-3-8B and Qwen3-8B, decode attention can contribute up to 68% of the total latency, as shown in Figure 1.
@@ 2.3. GPU Execution Model
Table 1. Memory hierarchy of A100-SXM4-80GB (nvidia2020a100; abdelkhalik2022demystifying).
@@
When executing an attention kernel, the workload is divided into CTAs, scheduled across SMs. During this process, the required KV cache data is transferred from slow global memory up the hierarchy into fast on-chip shared memory, where it is finally loaded into registers for computation. State-of-the-art implementations (dao2022flashattention; dao2023flashattention2; ye2025flashinfer) are designed to exploit this memory hierarchy. For example, FlashAttention (dao2022flashattention; nvidia2025cudabestpractices) partitions K/V caches into small tiles along the sequence length (Figure 3a) and processes them in a pipelined fashion: while one tile is computed, the next is prefetched asynchronously (Figure 3b), reducing memory latency.
@@ 3.1. Shared Prefixes and Prefix Reuse
Shared prefixes are prevalent in modern LLM workloads (openai2025promptcaching; wang2020generalizing; jin2024dvd), with studies showing that 40–62% of KV cache prefixes are shared across requests (qin2025mooncake; wang2025kvcache). These prefixes often form multi-level hierarchies (zheng2024sglang; pan2025fasttree). For example, a base system prompt shared by all users, followed by task-specific prefixes shared within subgroups.
Systems like SGLang(zheng2024sglang) and vLLM(kwon2023pagedattention) support prefix reuse by mapping shared logical prefixes to a single physical copy in GPU global memory. This reduces KV cache memory usage and improves concurrency. However, it can not reduce global memory access, which remains the decode attention bottleneck. To understand why, we analyze two key inefficiencies in current attention kernels under shared-prefix scenarios: redundant memory access (§3.2) and underutilized hardware resources (§3.3) as follows.
@@ 3.2. Redundant Memory Access
A major inefficiency in current attention kernels stems from redundant memory accesses. To execute a decode batch, attention kernels use a packing strategy that groups queries and their KV caches into CTAs. In the decode batch with 4 queries shown in Figure 4a, existing kernels’ query-centric paradigm adopts a one-query-per-CTA packing strategy (shah2024flashattention3; ye2025flashinfer), where each query and its KV blocks are independently assigned to a CTA (Figure 4b). While simple to schedule, this strategy causes shared KV prefixes (e.g., KV-0, KV-1) to be repeatedly loaded from slow global memory into on-chip memory. Although L2 cache offers partial reuse, it is limited by size and bandwidth.
@@ 3.3. Two-Dimension Resource Inefficiencies
Furthermore, existing attention kernels exhibit inefficient hardware utilization. Both query-centric and KV-centric designs use a tiled pipeline (Figure 3) to overlap memory access and computation. In this paradigm, selecting appropriate tile sizes (i.e., query tile size $m$ and KV tile size $n$ from Figure 3b) is critical for resource efficiency (dao2022flashattention). However, existing kernels adopt a one-size-fits-all design, which specifies a single, hard-coded tile size for all CTAs (e.g., $m=64,n=32$ (shah2024flashattention3)).
As shown in Figure 4b-c, this static approach ignores the dynamic nature of LLM workloads, leading to significant resource inefficiencies along two dimensions [^1] While specific prior works (pan2025fasttree) design one more query tile sizes $m$, it is still insufficient for handling highly dynamic query numbers and KV lengths. in Figure 5(b): (1) Memory Waste ($\mathbb{I}_{mem}$): When fewer than $m$ queries share a KV prefix, CTAs must pad inputs, wasting shared memory and registers to store unused data. (2) Execution Bubble ($\mathbb{I}_{exe}$): Due to varying KV lengths across CTAs, fixed-size tiling causes imbalanced workloads, leaving SMs underutilized in the tail stages of execution.
@@ 3.4. Insight: Pack-forward-merge Paradigm
Building on the observations above, we identify two design principles for efficient decode attention (Figure 4d): (1) Intra-CTA KV Cache Sharing: pack queries with shared KV prefixes into the same CTA to enable KV reuse in shared memory and avoid redundant global memory access. (2) Resource-Efficient Kernel Design: tailor kernel implementations to GPU architecture and CTA configurations to sustain high memory bandwidth and minimize resource inefficiencies.
@@ 4. Overview
To address these challenges, we design PAT, a memory-centric attention kernel implementation that follows the pack-forward-merge paradigm. It serves as a backend for the serving system vLLM (kwon2023pagedattention). In the pack stage, PAT adopts a profit-model-based heuristic packing strategy to aggregate queries that share KV into a CTA, so as to mitigate redundant global memory accesses (§5.1). It further designs multi-tile kernels and a runtime tile selector to choose an efficient kernel for each CTA (§5.2). In the forward stage, PAT adopts multi-stream execution and a long-KV split strategy to enable efficient kernel execution, which reduces execution bubbles (§6). Finally, in the merge stage, PAT uses a lightweight kernel based on online softmax to merge each query’s intermediate results across CTAs and produce the final output (§7).
@@ 5.1. Pack Scheduler
Problem Formulation Since decode attention is memory-bound (§2.3 and §3.2), the pack stage aims to minimize global memory accesses for a given decode batch (Figure 6a). The input is a batch of queries plus a block table where each row lists block IDs for a query; a shared prefix appears as identical leading block IDs across rows. The output is a partition $\mathcal{P}=\{P_{1},P_{2},\dots\}$ of CTAs, where each $P_{i}$ packs queries sharing successive identical prefix blocks (queries may be split across CTAs). We seek $\mathcal{P}^{\ast}$ that minimizes memory accesses, comprised of loaded KV cache size and per-query intermediate reads/writes due to splits/merges. The search space grows exponentially with query count and prefix lengths  (pan2025fasttree), so an exact solver is impractical for online serving, and we turn to a heuristic pack scheduler described next.
@@
In practice, the length of shared KV $l_{u}\geq 16$ since sharing is performed at the granularity of KV blocks (kwon2023pagedattention; shah2024flashattention3), whose sizes are typically larger than 16. Therefore, packing a node into a CTA yields a positive profit.
@@
Lazy Update. Although the pack scheduler is linear, its overhead can grow with batch size and number of KV blocks. We mitigate this with a lazy-update strategy that (1) reuses a scheduling result across continuous-batching iterations until the block table changes (e.g., request arrivals/departures or new KV block assignments) and (2) moves the scheduler into the serving system and runs it asynchronously once the block table is available. These reduce scheduler invocations from once per transformer layer to once several continuous-batching iterations and overlap scheduling with pre-attention task (e.g., LayerNorm, QKV projection), thereby substantially reducing the exposed scheduling latency (see §8.6) without affecting model accuracy.
@@ 5.2. Multi-tile Kernel
Insight and Approach. Given packed CTAs, PAT must choose per-CTA Q-tile $m$ and KV-tile $n$ to maximize resource efficiency. Because query sizes and KV lengths vary widely, single–tile-size kernels with padding (e.g., prior works (dao2023flashattention2; ye2025flashinfer; pan2025fasttree; zhu2024relayattention)) cannot adapt and suffer poor utilization (§3.3). To better exploit GPU shared memory and registers, we introduce a customized tiled attention kernel design that consists of: (1) a multi-tile kernel suite, where feasible $(m,n)$ configurations are derived from offline hardware and CTA-constraint analysis and implemented as resource-efficient kernels; and (2) a tile-size selector, an online decision-tree that selects the per-CTA $(m,n)$ to balance performance and parallelism.
@@ 1 $\mathcal{P}\leftarrow\emptyset$;
①Register and shared-memory constraints (upper bounds on $m,n$). Increasing $m$ or $n$ raises CTA shared-memory and register usage and can induce register spilling (nvidia2025cudaprogguide). To keep the kernel within hardware limits, we enforce two bounds: (1) Shared-memory constraint. One CTA’s shared-memory usage comprises the Q tile, K/V tile, and intermediate results (data type size $b^{\prime}$, usually higher precision). It must not exceed per-SM shared memory $S_{\mathrm{smem}}$:
@@
③CUTLASS constraint (lower bounds for $m$ and $n$): Efficient use of CUTLASS/CuTe MMA requires both tile sizes to be powers of two and at least [^5] It depends on data format; e.g., the minimum tile size is 32 for int8. 16 (cutlass_quickstart):
@@
Figure 8. Multi-tile kernel validation on H100-SXM4-80GB GPU. (a) Average bandwidth utilization under different tile sizes. (b) Average kernel latency under different tile sizes.:
Kernel equivalence. Per the offline constraint solving above, the candidate $(m,n)$ configurations sustain high bandwidth utilization, and decode-attention latency is dominated by global memory bandwidth. Thus, for decode batches without shared prefixes or execution bubbles, these configurations are performance-equivalent in bandwidth and latency. To validate this, we executed a decode batch without prefixes (KV length $1024$) under various configurations and used batch size $1134$, which is a common multiple of CTA concurrency across configurations on A100, avoiding execution bubbles. As shown in Figure 7(c) and Figure 7(d), all candidate configurations (underlined) sustain $83\%$-$86\%$ bandwidth utilization and exhibit similar end-to-end latency (difference $2\%$). This demonstrates that, in the absence of prefixes and bubbles, varying the tile configuration within the feasible set does not change CTA performance.
Porting PAT’s multi-tile kernel to other GPUs only requires re-deriving equivalent tile size configurations using constraints ①-③. On Hopper H100-SXM-80GB GPUs, this procedure removes the $(64,32)$ and $(64,64)$ configurations from Figure 7(b), and the remaining entries form the equivalent kernel set. Figure 8 reports validation on H100 at batch size $1188$, a common multiple of CTA concurrency across all configurations [^6] Batch size $1134$ for A100 and $1188$ for H100 are only used in the kernel-equivalence validation, not in the evaluation.. All equivalent configurations achieve $92.3\%$-$94.2\%$ bandwidth utilization and similar kernel latency. These results indicate that the same constraint-based procedure generalizes across architectures.